OpenStudy (anonymous):

Find the mean. 5.39, 8.35, 3.81

Parth (parthkohli):

(5.39 + 8.35 + 3.81) / 3 = 5.85

OpenStudy (callisto):

Mean =sum of data/ no. of data = (5.39+ 8.35+ 3.81)/3 = 17.55/3 =?

OpenStudy (anonymous):

5.39+8.35+3.81=17.55 17.55/3=5.85